# Improving Video Generation with Human Feedback

### v1

@@ Appendix F Hyperparameters
Training strategy LoRA (Hu et al., 2021)
@@
Optimizer Adam (Kingma, 2014)
@@
Table 10: Hyperparameters for Alignment Algorithms
@@
Optimizer Adam (Kingma, 2014)
@@
Optimizer Adam (Kingma, 2014)
@@
Table 11: Hyperparameters for reward modeling.
Appendix G Extended Experimental Results
Figure 10: Human evaluation of Flow-DPO on TA-Hard prompt.
Appendix H Prompt Subset of TA-Hard
A rabbit and a turtle racing on a track. The rabbit is sprinting ahead, while the turtle is steadily moving along. Spectators are cheering from the sidelines, and a finish line is visible in the distance.
A lion and a zebra playing soccer on a grassy field. The lion is dribbling the ball, while the zebra is trying to block it. The field is surrounded by trees, and other animals are watching the game.
A fox and an owl stargazing together on a hilltop. The fox is lying on its back, pointing at the stars, while the owl is perched on a nearby branch, looking through a telescope. The night sky is clear, with countless stars twinkling.
A dolphin and a sea turtle exploring a coral reef. The dolphin is swimming gracefully, while the sea turtle is gliding slowly beside it. The coral reef is vibrant with colorful corals and various marine life.
A dolphin and a whale singing together in the ocean. The dolphin is leaping out of the water, while the whale is producing deep, melodic sounds. The ocean is vast and blue, with the sun setting on the horizon.
A fox and a rabbit playing a duet on a piano in a forest clearing. The fox is playing the melody, while the rabbit is accompanying with harmony. The forest is alive with the sounds of nature, and other animals are gathered to listen.
A squirrel and a chipmunk building a treehouse in a large oak tree. The squirrel is hammering nails, while the chipmunk is holding a blueprint. The tree is tall and sturdy, with branches full of leaves.
A robot with glowing blue eyes and a human with a cybernetic arm playing basketball in a futuristic gym. The robot is dribbling the ball with precision, while the human is preparing to block the shot. The gym is equipped with advanced technology and holographic scoreboards.
A knight in shining armor and a wizard with a long, flowing beard practicing archery in a medieval courtyard. The knight is aiming at a target with a longbow, while the wizard is using magic to guide the arrows. The courtyard is surrounded by stone walls and blooming flowers.
A talking apple with eyes and a mouth, and a singing banana with legs hosting a talent show in a vibrant theater. The apple is the judge, giving feedback to contestants, while the banana is the host, entertaining the audience with jokes and songs. The theater is filled with colorful lights and excited spectators.
A pirate with a wooden leg and a mermaid with a shimmering tail playing a duet on a grand piano in an underwater cave. The pirate is playing the melody, while the mermaid is accompanying with harmony. The cave is illuminated by bioluminescent sea creatures, creating a magical atmosphere.
A superhero with a cape and a detective with a magnifying glass solving a mystery in a bustling city. The superhero is flying above the streets, scanning for clues, while the detective is examining evidence on the ground. The city is alive with activity, with skyscrapers towering overhead.
A chef with a tall hat and a robot with multiple arms cooking a gourmet meal in a state-of-the-art kitchen. The chef is chopping vegetables with precision, while the robot is simultaneously stirring, frying, and baking. The kitchen is equipped with the latest culinary technology, creating a seamless cooking experience.
A painter with a beret and a poet with a quill creating art in a sunlit studio. The painter is working on a vibrant canvas, while the poet is writing verses inspired by the artwork. The studio is filled with natural light and creative energy, with art supplies scattered around.
A spider with a square face and a green-furred puppy having a playful fight in a whimsical garden. The spider is using its web to swing around, while the puppy is playfully nipping at the spider’s legs. The garden is filled with oversized flowers and colorful mushrooms.
A talking teapot with a mustache and a dancing teacup with legs performing a tea ceremony in an enchanted forest. The teapot is pouring tea, while the teacup is twirling and dancing around. The forest is magical, with glowing plants and twinkling lights.
A robot with a television screen for a head and a toaster with arms and legs having a cooking competition in a retro kitchen. The robot is displaying recipes on its screen, while the toaster is popping out perfectly toasted bread. The kitchen is styled with vintage appliances and checkered floors.
A pair of animated scissors with eyes and a mouth and a roll of tape with tiny arms and legs wrapping presents in a festive workshop. The scissors are cutting wrapping paper with precision, while the tape is sealing the packages with a smile. The workshop is decorated with holiday lights and ornaments.
A pair of animated sneakers with eyes and a mouth and a talking basketball with a face playing a game of one-on-one on an urban basketball court. The sneakers are dribbling and making quick moves, while the basketball is bouncing and trying to score. The court is surrounded by graffiti-covered walls and cheering spectators.
A paper airplane with a scarf and a paper boat with a captain’s hat racing in the rain. The airplane glides through the air while the boat sails through puddles.
A basketball with a mohawk and a soccer ball with a bandana playing hopscotch in a playground. The basketball bounces high while the soccer ball rolls smoothly.
A mechanical knight with steam-powered joints standing guard at an ancient castle gate. Gears whir softly as its head turns to scan the surroundings, while steam occasionally escapes from its armor joints.
A wandering alchemist with potion-filled vials clinking on their belt, gathering herbs in an enchanted forest where mushrooms glow and flowers whisper secrets.
A mysterious plague doctor with clockwork enhancements peeking through their dark robes, mixing herbal remedies in a medieval apothecary shop as green smoke swirls from bubbling vials.

### v2

@@ Appendix I Hyperparameters
Training strategy LoRA [Hu et al., 2021]
@@
Optimizer Adam [Kingma, 2014]
@@
Table 11: Hyperparameters for reward modeling.
@@
Optimizer Adam [Kingma, 2014]
@@
Optimizer Adam [Kingma, 2014]
@@
GPUs 8 NVIDIA A800@80G
Appendix J Additional Qualitative Results
We present additional qualitative results generated by both the original model and the Flow-DPO aligned model, as shown in Fig.11 and Fig.12.
Figure 11: Additional visual comparison of videos generated by the original model and the Flow-DPO aligned model.
Figure 12: Additional visual comparison of videos generated by the original model and the Flow-DPO aligned model.
Appendix K Input Template for Reward Model
Full Input Template [VIDEO] You are tasked with evaluating a generated video based on three distinct criteria: Visual Quality, Motion Quality, and Text Alignment. Please provide a rating from 0 to 10 for each of the three categories, with 0 being the worst and 10 being the best. Each evaluation should be independent of the others. **Visual Quality:** Evaluate the overall visual quality of the video, with a focus on static factors. The following sub-dimensions should be considered: - **Reasonableness:** The video should not contain any significant biological or logical errors, such as abnormal body structures or nonsensical environmental setups. - **Clarity:** Evaluate the sharpness and visibility of the video. The image should be clear and easy to interpret, with no blurring or indistinct areas. - **Detail Richness:** Consider the level of detail in textures, materials, lighting, and other visual elements (e.g., hair, clothing, shadows). - **Aesthetic and Creativity:** Assess the artistic aspects of the video, including the color scheme, composition, atmosphere, depth of field, and the overall creative appeal. The scene should convey a sense of harmony and balance. - **Safety:** The video should not contain harmful or inappropriate content, such as political, violent, or adult material. If such content is present, the image quality and satisfaction score should be the lowest possible. Please provide the ratings of Visual Quality: <|VQ_reward|> END **Motion Quality:** Assess the dynamic aspects of the video, with a focus on dynamic factors. Consider the following sub-dimensions: - **Stability:** Evaluate the continuity and stability between frames. There should be no sudden, unnatural jumps, and the video should maintain stable attributes (e.g., no fluctuating colors, textures, or missing body parts). - **Naturalness:** The movement should align with physical laws and be realistic. For example, clothing should flow naturally with motion, and facial expressions should change appropriately (e.g., blinking, mouth movements). - **Aesthetic Quality:** The movement should be smooth and fluid. The transitions between different motions or camera angles should be seamless, and the overall dynamic feel should be visually pleasing. - **Fusion:** Ensure that elements in motion (e.g., edges of the subject, hair, clothing) blend naturally with the background, without obvious artifacts or the feeling of cut-and-paste effects. - **Clarity of Motion:** The video should be clear and smooth in motion. Pay attention to any areas where the video might have blurry or unsteady sections that hinder visual continuity. - **Amplitude:** If the video is largely static or has little movement, assign a low score for motion quality. Please provide the ratings of Motion Quality: <|MQ_reward|> END **Text Alignment:** Assess how well the video matches the textual prompt across the following sub-dimensions: - **Subject Relevance** Evaluate how accurately the subject(s) in the video (e.g., person, animal, object) align with the textual description. The subject should match the description in terms of number, appearance, and behavior. - **Motion Relevance:** Evaluate if the dynamic actions (e.g., gestures, posture, facial expressions like talking or blinking) align with the described prompt. The motion should match the prompt in terms of type, scale, and direction. - **Environment Relevance:** Assess whether the background and scene fit the prompt. This includes checking if real-world locations or scenes are accurately represented, though some stylistic adaptation is acceptable. - **Style Relevance:** If the prompt specifies a particular artistic or stylistic style, evaluate how well the video adheres to this style. - **Camera Movement Relevance:** Check if the camera movements (e.g., following the subject, focus shifts) are consistent with the expected behavior from the prompt. Textual prompt - [PROMPT] Please provide the ratings of Text Alignment: <|TA_reward|> END
Appendix L Prompt Subset of TA-Hard
A rabbit and a turtle racing on a track. The rabbit is sprinting ahead, while the turtle is steadily moving along. Spectators are cheering from the sidelines, and a finish line is visible in the distance. A lion and a zebra playing soccer on a grassy field. The lion is dribbling the ball, while the zebra is trying to block it. The field is surrounded by trees, and other animals are watching the game. A fox and an owl stargazing together on a hilltop. The fox is lying on its back, pointing at the stars, while the owl is perched on a nearby branch, looking through a telescope. The night sky is clear, with countless stars twinkling. A dolphin and a sea turtle exploring a coral reef. The dolphin is swimming gracefully, while the sea turtle is gliding slowly beside it. The coral reef is vibrant with colorful corals and various marine life. A dolphin and a whale singing together in the ocean. The dolphin is leaping out of the water, while the whale is producing deep, melodic sounds. The ocean is vast and blue, with the sun setting on the horizon. A fox and a rabbit playing a duet on a piano in a forest clearing. The fox is playing the melody, while the rabbit is accompanying with harmony. The forest is alive with the sounds of nature, and other animals are gathered to listen. A squirrel and a chipmunk building a treehouse in a large oak tree. The squirrel is hammering nails, while the chipmunk is holding a blueprint. The tree is tall and sturdy, with branches full of leaves. A robot with glowing blue eyes and a human with a cybernetic arm playing basketball in a futuristic gym. The robot is dribbling the ball with precision, while the human is preparing to block the shot. The gym is equipped with advanced technology and holographic scoreboards. A knight in shining armor and a wizard with a long, flowing beard practicing archery in a medieval courtyard. The knight is aiming at a target with a longbow, while the wizard is using magic to guide the arrows. The courtyard is surrounded by stone walls and blooming flowers. A talking apple with eyes and a mouth, and a singing banana with legs hosting a talent show in a vibrant theater. The apple is the judge, giving feedback to contestants, while the banana is the host, entertaining the audience with jokes and songs. The theater is filled with colorful lights and excited spectators. A pirate with a wooden leg and a mermaid with a shimmering tail playing a duet on a grand piano in an underwater cave. The pirate is playing the melody, while the mermaid is accompanying with harmony. The cave is illuminated by bioluminescent sea creatures, creating a magical atmosphere. A superhero with a cape and a detective with a magnifying glass solving a mystery in a bustling city. The superhero is flying above the streets, scanning for clues, while the detective is examining evidence on the ground. The city is alive with activity, with skyscrapers towering overhead. A chef with a tall hat and a robot with multiple arms cooking a gourmet meal in a state-of-the-art kitchen. The chef is chopping vegetables with precision, while the robot is simultaneously stirring, frying, and baking. The kitchen is equipped with the latest culinary technology, creating a seamless cooking experience. A painter with a beret and a poet with a quill creating art in a sunlit studio. The painter is working on a vibrant canvas, while the poet is writing verses inspired by the artwork. The studio is filled with natural light and creative energy, with art supplies scattered around. A spider with a square face and a green-furred puppy having a playful fight in a whimsical garden. The spider is using its web to swing around, while the puppy is playfully nipping at the spider’s legs. The garden is filled with oversized flowers and colorful mushrooms. A talking teapot with a mustache and a dancing teacup with legs performing a tea ceremony in an enchanted forest. The teapot is pouring tea, while the teacup is twirling and dancing around. The forest is magical, with glowing plants and twinkling lights. A robot with a television screen for a head and a toaster with arms and legs having a cooking competition in a retro kitchen. The robot is displaying recipes on its screen, while the toaster is popping out perfectly toasted bread. The kitchen is styled with vintage appliances and checkered floors. A pair of animated scissors with eyes and a mouth and a roll of tape with tiny arms and legs wrapping presents in a festive workshop. The scissors are cutting wrapping paper with precision, while the tape is sealing the packages with a smile. The workshop is decorated with holiday lights and ornaments. A pair of animated sneakers with eyes and a mouth and a talking basketball with a face playing a game of one-on-one on an urban basketball court. The sneakers are dribbling and making quick moves, while the basketball is bouncing and trying to score. The court is surrounded by graffiti-covered walls and cheering spectators. A paper airplane with a scarf and a paper boat with a captain’s hat racing in the rain. The airplane glides through the air while the boat sails through puddles. A basketball with a mohawk and a soccer ball with a bandana playing hopscotch in a playground. The basketball bounces high while the soccer ball rolls smoothly. A mechanical knight with steam-powered joints standing guard at an ancient castle gate. Gears whir softly as its head turns to scan the surroundings, while steam occasionally escapes from its armor joints. A wandering alchemist with potion-filled vials clinking on their belt, gathering herbs in an enchanted forest where mushrooms glow and flowers whisper secrets. A mysterious plague doctor with clockwork enhancements peeking through their dark robes, mixing herbal remedies in a medieval apothecary shop as green smoke swirls from bubbling vials.